A quantity of 7 units total costs $194 and a quantity of 8 units total costs is $208. What is the marginal cost for 8 units? How did you arrive to this solution? How do you define marginal cost?
Answer #1

Answer

Marginal Cost = Change in in Total Cost / Change in Quantity

When Quantity = 8 , Total cost = 208 and When Quantity = 7 , Total cost = 194

=> Marginal Cost = (208 - 194)/(8 - 7) = 14

Hence, the marginal cost for 8 units is $14.

Marginal Cost is defined as the additional cost incurred in order to produce one additional unit of output.

Note:

We can also say that MC(n) = TC(n) - TC(n - 1)

where MC(n) = Marginal cost of producing nth unit and TC(n) = Total cost of producing n units

the marginal cost for 8 units = MC(*) = TC(8) - TC(7) = 208 - 194 = $14